# Fanxu Meng1,2,3, Zhaohui Wang1, Muhan Zhang1,3

### v2

@@ Abstract
As the parameters of large language models (LLMs) expand, the computational cost of fine-tuning the entire model becomes prohibitive. To address this challenge, we introduce a parameter-efficient fine-tuning (PEFT) method, Principal Singular values and Singular vectors Adaptation (PiSSA), which optimizes a significantly reduced parameter space while achieving or surpassing the performance of full-parameter fine-tuning. PiSSA is inspired by Intrinsic SAID, which suggests that pre-trained, over-parametrized models inhabit a space of low intrinsic dimension. Consequently, PiSSA represents a matrix $W\in\mathbb{R}^{m\times n}$ within the model by the product of two trainable matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$, plus a residual matrix $W^{res}\in\mathbb{R}^{m\times n}$ for error correction. Singular value decomposition (SVD) is employed to factorize $W$, and the principal singular values and vectors of $W$ are utilized to initialize $A$ and $B$. The residual singular values and vectors initialize the residual matrix $W^{res}$, which keeps frozen during fine-tuning. Notably, PiSSA shares the same architecture with Low-Rank Adaptation (LoRA), which hypothesizes that changes in model parameters $\Delta W$ form a low-rank matrix. However, LoRA approximates $\Delta W$ through the product of two matrices, $A$, initialized with Gaussian noise, and $B$, initialized with zeros, while PiSSA initializes $A$ and $B$ with principal singular values and singular vectors of the original matrix $W$. Given that the principal singular values and vectors capture the essence of a low-rank matrix, PiSSA can better approximate the outcomes of full-parameter fine-tuning at the beginning by changing the essential parts while freezing the “noisy” parts. In comparison, LoRA freezes the original matrix and updates the “noise”. This distinction enables PiSSA to convergence much faster than LoRA and also achieve better performance in the end. On five common benchmarks, PiSSA outperforms LoRA on all of them using exactly the same setups except for a different initialization. On GSM8K, Mistral-7B fine-tuned with PiSSA achieves an accuracy of 72.86%, outperforming LoRA’s 67.7% by 5.16%. Due to the same architecture, PiSSA inherits many of LoRA’s advantages, such as parameter efficiency and compatibility with quantization. Furthermore, PiSSA reduces the 4-bit quantization error in LLaMA 2-7B by 18.97%, resulting in a substantial improvement in fine-tuning performance. On the GSM8K benchmark, PiSSA achieves an accuracy of 49.13%, surpassing the performances of QLoRA at 39.8% and LoftQ at 40.71%. Leveraging a fast SVD technique, the initialization of PiSSA takes only a few seconds, inducing negligible cost of switching LoRA to PiSSA.
1 Introduction
LoRA [11] hypothesizes that the modifications to parameter matrices during fine-tuning exhibit low-rank properties. As depicted in Figure 1(b), for a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA substitutes the updates with a low-rank decomposition $\Delta W=AB$, where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, and the rank $r\ll\text{min}(m,n)$. For $h=Wx$, the modified forward pass computation is expressed as follows:
$h=(W+\Delta W)x=(W+AB)x.$ (1)
A random Gaussian initialization is used for $A$ and zero for $B$, making $AB=0$ at the beginning of training, thereby the injection of adapters does not affect the model’s output initially. LoRA avoids the need to compute gradients or maintain the optimizer states for the original matrix, instead optimizing the injected, significantly smaller low-rank matrices. Thus, it could reduce the number of trainable parameters by 10,000 times and the GPU memory requirement by 3 times. Moreover, LoRA often achieves comparable or superior performance to full parameter fine-tuning, indicating that fine-tuning “parts” of the full parameters can be enough for downstream tasks. By integrating the quantization of pre-trained models, LoRA also enables reducing the average memory requirements by 16 times [10]. Meanwhile, the adapters are still allowed to utilize higher precision weights; thus, the quantization usually does not degrade the performance of full-precision base models.
@@ 1 Introduction
Figure 1: The comparison among Full Fine-tuning, training with LoRA, and PiSSA. In this visualization, blue modules represent parts of the model where parameters are frozen during training, while orange modules indicate components that require updates.
Inspired by the core idea of Intrinsic SAID [14, 15], our Principal Singular values and Singular vectors Adapter (PiSSA) diverges from LoRA and its successors by not focusing on approximating $\Delta W$, but $W$. In alignment with Intrinsic SAID, we assume that the matrix $W$ in pre-trained models exhibits low (intrinsic) rank characteristics. Thus, we apply singular value decomposition (SVD) to matrix $W$. Based on the magnitude of the singular values, we divide $W$ into two components: the principal matrix $W^{pri}$ and the residual matrix $W^{res}$. The principal matrix $W^{pri}$ is of low rank and can be represented by the product of $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$. $A$ and $B$ are initialized based on the principal singular values and vectors and are trainable, whereas $W^{res}$ is initialized with the product of the residual singular values and vectors and are frozen. As indicated by Equation (2), the integration of $AB$ (the PiSSA adaptor) with the residual model also preserves the capability of the pre-trained model in the beginning of fine-tuning:
$h=Wx=(W^{res}+W^{pri})x=(W^{res}+AB)x.$ (2)
Different from LoRA, a significantly less important part $W^{res}$ is kept frozen, while the essential part $W^{pri}$ is fully tunable from the beginning, which enables PiSSA to fit the training data faster and better. Since the high-frequency principal component $W^{pri}$ is preserved in the adapter, when fine-tuning PiSSA incorporated with quantization, PiSSA significantly reduced quantization errors in the initial stages compared to using QLoRA.
@@ 2 Related Works
The vast complexity and computational needs of large language models (LLMs) with billions of parameters present significant hurdles in adapting them for specific downstream tasks. Parameter Efficient Fine-Tuning (PEFT) [12, 13] emerges as a compelling solution by minimizing the fine-tuning parameters and memory requirements while achieving comparable performance to full fine-tuning. PEFT encompasses strategies like partial fine-tuning, soft prompt fine-tuning, non-linear adapter fine-tuning, and linear adapter fine-tuning.
Partial fine-tuning selects certain model parts, e.g., bias and layer normalization [16, 17], or inserted trainable mask to the pre-trained weights [18, 19] and delta weights [20, 21, 22, 23], for fine-tuning. However, the performance of partial fine-tuning may be limited due to only specific layers’ parameters being fine-tuned. Soft prompt fine-tuning  [24, 25, 26, 27, 28, 29, 30] introduces optimizable vectors into the model’s input or hidden layers, adapting to specific tasks with minimal structural changes. Non-linear adapter fine-tuning  [31, 32, 33, 34, 35, 36, 37, 38, 39, 40] incorporates (in a non-linear manner) trainable (non-linear) modules to the models. Both soft prompt fine-tuning and non-linear adapter fine-tuning add complexity and inference costs despite their adaptability. Linear adapter fine-tuning [14, 15, 11, 41], particularly LoRA [11] and its derivative methods, involves adding trainable parameters to the self-attention and feed-forward layers, which are then linearly combined with the original weights. After fine-tuning, these adaptations can be reparametrized into the standard model structure, thus gaining widespread adoption due to their ability to maintain the model’s original architecture while enabling efficient fine-tuning.
LoRA [11] posits that weight adjustments during model adaptation exhibit a low "intrinsic rank". It learns the weight updates through a low-rank decomposition, keeping the original structure frozen and proving effective in parameter-efficient fine-tuning. Following LoRA, Low-rank Adjustment [42, 41, 43] learns the rank size needed for LoRA in each layer of the model. LoRA-guided Pretrained Weight Update [44, 45] updates the original weights of the model using the parameters from adapter layers, enhancing LoRA’s representational capacity. Quantization Adaptation [10, 46, 47] applies quantization to the base model, further reducing training consumption overhead.
Differing from LoRA’s approach of learning low-rank approximations of weight updates, our PiSSA method innovates by utilizing the low intrinsic dimension of pre-trained weight directly. It initializes updatable adapter layers with principal singular values and vectors from the model’s weights, preserving the essential capabilities while enabling effective fine-tuning. This approach, while different from LoRA’s in philosophy, shares most of the structural benefits of LoRA and can be combined with methods like quantization the residual model for reducing computation cost.
3 Method
@@
$A$ $=U_{[:r]}\,\text{diag}(S_{[:r]}^{1/2})\in\mathbb{R}^{m\times r},$ (3)
$B$ $=\text{diag}(S_{[:r]}^{1/2})\,V_{[:r]}^{T}\in\mathbb{R}^{r\times n}.$ (4)
@@
The residual singular values and vectors are used to build the residual matrix:
$W^{res}=U_{[r:]}\,\text{diag}(S_{[r:]})\,V_{[r:]}^{T}\in\mathbb{R}^{m\times n}.$ (5)
@@
Given $W=W^{res}+AB$, the introduction of PiSSA in the beginning of training does not compromise the model’s capabilities. Considering $S_{[:r]}\gg S_{[r:]}$, thus rendering $W^{pri}\gg W^{res}$ and $W\approx AB$, the adapter $AB$ encapsulates the pre-trained model’s primary capabilities. Then, the fine-tuning of $AB$ mirrors the process of fine-tuning the entire model, despite using much less parameters. In contrast, LoRA initializes the adapter with Gaussian noise and zeros, thus the principal part of $W$ is not updatable in the early stage of the fine-tuning, possibly introducing much waste of gradient descent steps. Moreover, an inferior initialization might lead to suboptimal local minimum points found, causing worse generalization performance.
@@
Since PiSSA shares the identical architecture with LoRA, it inherits most of LoRA’s benefits. These include, but are not limited to, the capability of fine-tuning a model with a reduced number of trainable parameters, quantizing the residual model to decrease memory consumption during forward propagation in training, and easy deployment, during which process PiSSA can significantly reduce the initial quantization error in the initial stages. The adapter’s straightforward linear structure facilitates the integration of trainable matrices with the pre-trained weights upon deployment, thereby maintaining the original inference speed of a fully fine-tuned model. For storage efficiency, we can choose not to store the dense parameter matrix $\Delta W$, but to store the low-rank matrices, $\Delta AB$ instead. As shown in Appendix A, leveraging solely the $\Delta AB$ facilitates their seamless integration with the original pre-trained models. Finally, one pre-trained model can accommodate multiple $\Delta AB$, fine-tuned by diverse PiSSA procedures, which enables fast adapting of the pre-trained model to different downstream applications.

### v3

@@ Abstract
To parameter-efficiently fine-tune (PEFT) large language models (LLMs), the low-rank adaptation (LoRA) method approximates the model changes $\Delta W\in\mathbb{R}^{m\times n}$ through the product of two matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$, $A$ is initialized with Gaussian noise, and $B$ with zeros. LoRA freezes the original model $W$ and updates the “Noise & Zero” adapter, which may lead to slow convergence. To overcome this limitation, we introduce Principal Singular values and Singular vectors Adaptation (PiSSA). PiSSA shares the same architecture as LoRA, but initializes the adaptor matrices $A$ and $B$ with the principal components of the original matrix $W$, and put the remaining components into a residual matrix $W^{res}\in\mathbb{R}^{m\times n}$ which is frozen during fine-tuning. Compared to LoRA, PiSSA updates the principal components while freezing the “residual” parts, allowing faster convergence and enhanced performance. Comparative experiments of PiSSA and LoRA across 12 different models, ranging from 184M to 70B, encompassing 5 NLG and 8 NLU tasks, reveal that PiSSA consistently outperforms LoRA under identical experimental setups. On the GSM8K benchmark, Mistral-7B fine-tuned with PiSSA achieves an accuracy of 72.86%, surpassing LoRA’s 67.7% by 5.16%. Due to the same architecture, PiSSA is also compatible with quantization to further reduce the memory requirement of fine-tuning. Compared to QLoRA, QPiSSA (PiSSA with 4-bit quantization) exhibits smaller quantization errors in the initial stages. Fine-tuning LLaMA-3-70B on GSM8K, QPiSSA attains an accuracy of 86.05%, exceeding the performances of QLoRA at 81.73%. Leveraging a fast SVD technique, PiSSA can be initialized in only a few seconds, presenting a negligible cost for transitioning from LoRA to PiSSA.
@@
Figure 1: The comparison among Full Fine-tuning, training with LoRA, and PiSSA. In this visualization, blue modules represent parts of the model where parameters are frozen during training, while orange modules indicate components that require updates. QLoRA quantizes the pretrained matrix in LoRA to 4-bit, whereas QPiSSA quantizes the residual matrix in PiSSA.
(a) PiSSA converges more rapidly.
(b) PiSSA reduces more quantization error.
Figure 2: We illustrate the two key advantages of PiSSA: converging faster and better, and reducing quantization error. In the left figure, we use a toy example to show PiSSA’s faster convergence, where we first train a two-layer MLP classifying odd numbers of MNIST, and then fine-tune the model on even numbers. PiSSA finds the right direction more quickly and achieves a lower loss with the same number of steps. In the right figure, we show the quantization error reduction of PiSSA compared to directly quantizing the base model. PiSSA can reduce more quantization error than the baselines.
1 Introduction
LoRA [11] hypothesizes that the modifications to parameter matrices during fine-tuning exhibit low-rank properties. As depicted in Figure 1(b), for a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA substitutes the updates with a low-rank decomposition $\Delta W=AB$, where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, and the rank $r\ll\text{min}(m,n)$. For $Y=XW$, the modified forward pass is as follows:
$Y=X(W+\Delta W)=X(W+AB),$ (1)
where $X\in\mathbb{R}^{b\times m}$, $Y\in\mathbb{R}^{b\times n}$, and $b$ represents the batch size of input data. A random Gaussian initialization is used for $A$ and zero for $B$, making $AB=0$ at the beginning of training, thereby the injection of adapters does not affect the model’s output initially. LoRA avoids the need to compute gradients or maintain the optimizer states for the original matrix $W$, instead optimizing the injected, significantly smaller low-rank matrices $A,B$. Thus, it could reduce the number of trainable parameters by 10,000 times and the GPU memory requirement by 3 times [11]. Moreover, LoRA often achieves comparable or superior performance to full parameter fine-tuning, indicating that fine-tuning “parts” of the full parameters can be enough for downstream tasks. By integrating the quantization of pre-trained matrices $W$, LoRA also enables reducing the average memory requirements by 16 times [10]. Meanwhile, the adapters are still allowed to utilize higher precision weights; thus, the quantization usually does not significantly degrade the performance of LoRA.
According to Equation 1, the gradients of A and B are $\frac{\partial L}{\partial A}=X^{T}\left(\frac{\partial L}{\partial Y}\right)B^{T}$ and $\frac{\partial L}{\partial B}=A^{T}X^{T}\left(\frac{\partial L}{\partial Y}\right)$. Compared to full fine-tuning, using LoRA initially does not change the output $Y$ for the same input $X$, so the gradient magnitude is primarily determined by the values of $A$ and $B$. Since $A$ and $B$ are initialized with Gaussian noise and zeros in LoRA, the gradients can be very small, leading to slow convergence in the fine-tuning process. We also observe this phenomenon empirically, where LoRA often wastes much time around the initial point.
Our Principal Singular values and Singular vectors Adapter (PiSSA) diverges from LoRA and its successors by focusing not on approximating $\Delta W$, but $W$. We apply singular value decomposition (SVD) to matrix $W$. Based on the magnitude of the singular values, we partition $W$ into two parts: the principal low-rank matrix $W^{pri}$, comprising a few largest singular values, and the residual matrix $W^{res}$, which possesses the remaining smaller singular values (with a larger quantity, representing a possible long-tail distribution). The principal matrix $W^{pri}$ can be represented by the product of $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$. As depicted in Figure 1(c), $A$ and $B$ are initialized based on the principal singular values and singular vectors and are trainable. Conversely, $W^{res}$ is initialized with the product of the residual singular values and singular vectors and remains frozen during fine-tuning. Since the principal singular vectors represent the directions in which the matrix $W$ has the most significant stretching or impact, by directly tuning these principal components, PiSSA is able to fit the training data faster and better (as demonstrated in Figure 2(a)). Moreover, the loss and gradient norm curves of PiSSA often demonstrate a similar trend to those of full parameter fine-tuning in our experiments (Figure 4), indicating that fine-tuning the principal components matches the behavior of fine-tuning the full matrix to some degree.
Since the principal components $W^{pri}$ are preserved in the adapter and are of full precision, another benefit of PiSSA is that when applying quantization to the frozen part $W^{res}$, we can significantly reduce the quantization error compared to QLoRA (which quantizes the whole $W$), as illustrated in Figure 2(b). Therefore, PiSSA is perfectly compatible with quantization, making it a plug-and-play substitution for LoRA.
@@ 2 Related Works
The vast complexity and computational needs of large language models (LLMs) with billions of parameters present significant hurdles in adapting them for specific downstream tasks. Parameter Efficient Fine-Tuning (PEFT) [12, 13] emerges as a compelling solution by minimizing the fine-tuning parameters and memory requirements while achieving comparable performance to full fine-tuning. PEFT encompasses strategies like partial fine-tuning [14, 15, 16, 17, 18, 19, 20, 21], soft prompt fine-tuning  [22, 23, 24, 25, 26, 27, 28], non-linear adapter fine-tuning [29, 30, 31, 32, 33, 34, 35, 36, 37, 38], and low rank adapter based fine-tuning [39, 40, 11, 41].
LoRA [11] inject trainable adapters to the linear layers. After fine-tuning, these adaptations can be re-parameterized into the standard model structure, thus gaining widespread adoption due to their ability to maintain the model’s original architecture while enabling efficient fine-tuning. Following LoRA, AdaLoRA [42, 41, 43] dynamically learns the rank size needed for LoRA in each layer of the model. DeltaLoRA [44, 45] updates the original weights of the model using parameters from adapter layers, enhancing LoRA’s representational capacity. LoSparse [46] incorporates LoRA to prevent pruning from eliminating too many expressive neurons. DoRA [47] introduces a magnitude component to learn the scale of $\Delta W$ while utilizing the original AB as a direction component of $\Delta W$. Unlike LoRA and its successors, which focus on learning low-rank approximations of weight updates, our PiSSA approach directly tunes the essential but low-rank parts of the model while keeping the noisier, high-rank, and nonessential parts frozen. Although our approach differs in philosophy from LoRA, it shares most of LoRA’s structural benefits and can be extended by these methods to enhance its performance.
QLoRA [10] integrates LoRA with 4-bit NormalFloat (NF4) quantization, along with Double Quantization and Paged Optimizers, enabling the fine-tuning of a 65B parameter model on a single 48GB GPU while preserving the performance of full 16-bit fine-tuning tasks. QA-LoRA [48] introduces group-wise operators to increase the degree of freedom in low-bit quantization. LoftQ [49] reduces quantization error by decomposing the quantization error matrix of QLoRA and retaining the principal components with an adapter. Our PiSSA approach can also be combined with quantization techniques, and we have found that PiSSA significantly reduces quantization error compared to QLoRA and LoftQ.
3 PiSSA: Principal Singular Values and Singular Vectors Adaptation
@@
$A$ $=U_{[:,:r]}\,S_{[:r,:r]}^{1/2}\in\mathbb{R}^{m\times r},$ (2)
$B$ $=S_{[:r,:r]}^{1/2}\,V_{[:,:r]}^{T}\in\mathbb{R}^{r\times n}.$ (3)
@@
The residual singular values and vectors are used to build the residual matrix which is frozen during fine-tuning:
$W^{res}=U_{[:,r:]}\,S_{[r:,r:]}\,V_{[:,r:]}^{T}\in\mathbb{R}^{m\times n}.$ (4)
As indicated by Equation 5, the integration of $AB$ with the residual matrix also preserves the full capability of the pre-trained model in the beginning of fine-tuning:
$Y=XW=X(W^{res}+W^{pri})=X(W^{res}+AB).$ (5)
@@
Similar to LoRA, the gradients of $A$ and $B$ are also given by $\frac{\partial L}{\partial A}=X^{T}\left(\frac{\partial L}{\partial Y}\right)B^{T}$ and $\frac{\partial L}{\partial B}=A^{T}X^{T}\left(\frac{\partial L}{\partial Y}\right)$. Since elements of $\mathbf{s}_{[:r]}$ $\gg$ elements of $\mathbf{s}_{[r:]}$, the trainable adapter $W^{pri}=AB$ contains the most essential directions of $W$. In the ideal case, training $AB$ mirrors the process of fine-tuning the entire model despite using fewer parameters. The ability to directly fine-tune the most essential part of a model enables PiSSA to converge faster and better. In contrast, LoRA initializes the adapters $A$ and $B$ with Gaussian noise and zeros while keeping $W$ frozen. Consequently, the gradients are small or in random directions during the early stages of fine-tuning, possibly introducing much waste of gradient descent steps. Moreover, an inferior initialization might lead to suboptimal local minimum points found, causing worse generalization performance.
@@
Since PiSSA shares the identical architecture with LoRA, it inherits most of LoRA’s benefits. These include but are not limited to the capability of fine-tuning a model with a reduced number of trainable parameters, quantizing the residual model to decrease memory consumption during forward propagation in training, and easy deployment. The adapter’s straightforward linear structure facilitates the integration of trainable matrices with the pre-trained weights upon deployment, thereby maintaining the original inference speed of a fully fine-tuned model. Employing the Fast SVD technique [50] allowed PiSSA to finish initialization in several seconds (Appendix B), which is a negligible cost.
For storage efficiency, we can choose not to store the dense parameter matrix $\Delta W$, but to store the low-rank matrices, $\Delta A$ and $\Delta B$ instead. As shown in Appendix C, leveraging solely the $\Delta A$ and $\Delta B$ facilitates their seamless integration with the original pre-trained models. Finally, one pre-trained model can accommodate multiple $\Delta A,\Delta B$, fine-tuned by diverse PiSSA or LoRA procedures, which enables fast adaptation of the pre-trained model to different downstream applications.
4 QPiSSA: PiSSA with Quantization
Quantization divides the value range of a matrix into several continuous regions, and maps all values falling inside a region into the same “quantized” value. It is an effective technique to reduce the memory consumption of forward propagation, but breaks down during backpropagation. At the same time, LoRA greatly reduces the backward memory requirement, making it highly suitable to use LoRA and quantization together, where the base model is quantized for memory-efficient forward propagation, and the LoRA adaptors are kept in full precision for accurate backward parameter updates. One representative previous work, QLoRA, quantizes the base model to Normal Float 4-bit (NF4) and initializes the full-precision $A$ and $B$ with Gaussian-Zero initialization. Therefore, the overall error is given by:
Quantization Error of $\text{Quantization Error of QLoRA}=||W-\left(nf4(W)+AB\right)||_{*}=||W-nf4(W)||_{*},$ (6)
where $||M||_{*}$ denotes the nuclear norm (also known as the trace norm)  [51], defined as:
$\|M\|_{*}=\operatorname{trace}\left({\sqrt{M^{*}M}}\right)=\sum_{i=1}^{\min\{m,n\}}\sigma_{i}(M),$ (7)
where $\sigma_{i}(M)$ is the $i^{\text{th}}$ singular value of $M$. As we can see, the quantization error of QLoRA is the same as that of directly quantizing the base model. Our QPiSSA, however, does not quantize the base model but the residual model. Therefore, its error is given by:
Quantization Error of $\text{Quantization Error of QPiSSA}=||W-\left(nf4(W^{res})+AB\right)||_{*}=||W^{res}-nf4(W^{res})||_{*}.$ (8)
Since the residual model has removed the large-singular-value components, $W^{res}$ has a narrower distribution than that of $W$, as can be seen in Figures 3(a) and 3(b) (comparing the singular value distributions of $W$ and $W^{res}$), as well as Figures 3(c) and 3(f) (comparing the value distributions of $W$ and $W^{res}$), which is beneficial for reducing the quantization error. Furthermore, since NF4 is optimized for normally distributed data [10], we fit a Gaussian for the values of $W$ and $W^{res}$, respectively. As can be seen in Figures 3(c) and 3(f), $W^{res}$ is much more Gaussian-like and has a smaller standard deviation, making it more suitable to apply NF4 to $W^{res}$ instead of $W$. Both the above lead QPiSSA to achieve a significantly lower quantization error than QLoRA, shown in Figures 3(d) and 3(e).
(a) Original matrix $W$
(b) Residual matrix $W^{res}$
(c) The distribution of $W$
(d) Error matrix of QLoRA
(e) Error matrix of QPiSSA
(f) The distribution of $W^{res}$
Figure 3: Figures (a), (b), (d) and (e) visualize the singular values in descending order for $W$, $W^{res}$, $W-nf4(W)$ and $W^{res}-nf4(W^{res})$ of LLaMA 2-7B’s “layers[0].self_attn.q_proj” matrix. Figures (c) and (f) show the data value distributions of $W$ and $W^{res}$.
Besides the advantage of reducing quantization error, QPiSSA’s gradient direction is similar to that of PiSSA, resulting in significantly better fine-tuning performance compared to QLoRA.